How to solve this word problem?
2 cars started off in the same direction at the same time, but the first car was going twice as fast as the second car. In 5 hrs, the first car was ahead of the second car by 275 miles. How fast was going the second car?

Answer by checkley77(12569)

5x=5(2x)-275
5x=10x-275
5x-10x=275
-5x=-275
x=-275/-5
x=55 mph for the slower/second car.
proof:
5*55=10*55=-275
275=55-275
275=275